During the time when a man had flu, he ran a fever of 2.0oC above normal. His body temperature was 39.0oC instead of the normal 37.0oC. Assuming that the man has a mass of 80 kg and that the human body is mostly water, how much heat is required to raise his temperature? [Take specific heat capacity of liquid as, C = 4186 J / kg.oC]


1
Expert's answer
2021-05-03T10:27:54-0400

"Q=cm\Delta{T}=4186\frac{J}{kg\cdot\degree{C}}\times80kg\times2.0\degree{C}=6.7\times10^5J=670kJ"


Answer: 670 kJ